OpenStudy (hulkart):

HELP FAST!!!!!! Given the relation y = 3x − 2, identify the x- and y- axis intercepts.

OpenStudy (shamim):

Put x=0 in ur given equation . Tell me y=?

OpenStudy (shamim):

The value of y is called y intercept

OpenStudy (hulkart):

uhh...

OpenStudy (hockeychick23):

y = 3x − 2 0= 3x-2 0+2=3x 2/2=3/2x x=2/3 y=3(0)-2 y= -2

OpenStudy (sepeario):

Basically what he's saying is to find the x intercept, you make y zero and solve for x. The same may apply for the y intercept, but you can just find that by the linear equation y=mx+c.

OpenStudy (anonymous):

As it says "identify" instead of find, i think it means what Sepeario says. Just recognize the value of c and m in the equation....